OpenStudy (anonymous):

The ratio of two numbers is 5:7. If 6 is added to each number, the ratio of the numbers is 7:9. What are the numbers?

OpenStudy (anonymous):

difference between the numbers is 2. 6/2 = 3 3 x 5 = 15 3 x 7 = 21 numbers are 15 and 21

OpenStudy (anonymous):

How did you find the difference of the numbers equals 2???

OpenStudy (anonymous):

5 : 7 +2 +2 7 : 9

OpenStudy (anonymous):

I get that,but 6 is added to each not 2???

OpenStudy (anonymous):

yes but these are scaled down so we need to find how much they are scaled down to, by finding the ratio to the two ratios we can find that they are scaled down by 3.

OpenStudy (anonymous):

I'm sorry,but I don't quit understand.

OpenStudy (anonymous):

I know its hard to explain over chat... hmm.. how can I put this.

OpenStudy (anonymous):

The part I don't understand is the scaled down part.

OpenStudy (anonymous):

alright well we know that the numbers are not 5 and 7, these are just simplified versions of the real numbers.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so that means that 5:7 to 7:9 just adds 2 to both the numbers, not 6 lie the problem says. so how do we get from 2 to 6?

OpenStudy (anonymous):

I'm lost on how to get from 2 to 6.

OpenStudy (anonymous):

we always multiply when we are talking ratios..

OpenStudy (anonymous):

2 x 3 = 6

OpenStudy (anonymous):

so 3 is our scale for the whole entire equation

OpenStudy (anonymous):

lets multiply all our numbers by our scale 5:7 7:9 15:21 21:28

OpenStudy (anonymous):

What do you mean by scale? i guess i haven't learned this term yet

OpenStudy (anonymous):

Usually I work these problems by cross multiplying if that is any help.

OpenStudy (anonymous):

Ill find out how to explain this indeed! :)

OpenStudy (anonymous):

these are simplified ratios. which means that they were simplified with one single number correct?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

we are trying to find that one number. we are given that between the two numbers, 6 is added, but since it is scaled down, so is that 6.

OpenStudy (anonymous):

hold on... this might take a minute for me to understand.

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

i am sorry. i am lost again

OpenStudy (anonymous):

example: 3:6 + 6 = 9:12

OpenStudy (anonymous):

ok doky

OpenStudy (anonymous):

this WHOLE equation can be simplified by one number.. do you know what that number is?

OpenStudy (anonymous):

6??? or 2???

OpenStudy (anonymous):

one number that goes into 3, 6, 9 and 12

OpenStudy (anonymous):

what???

OpenStudy (anonymous):

3 goes into all those numbers

OpenStudy (anonymous):

so we simplify our equation and divide 3 by all the numbers. 1:3 + 2 = 3:4 Confused there?

OpenStudy (anonymous):

Oh duuu!!! I am sorry my brain isn't cooperating right now.:)

OpenStudy (anonymous):

lol its all good :) so the number that we added to the first ratio changed from 6 to 2. We are trying to work this backwards, this is what we have. 5:7 + 2 = 7:9 x:x + 6 = x:x what number can we mulitpy everything by to get the top equation to look like the bottom?

OpenStudy (anonymous):

ignoring the x's and focusing on the number we do have, 6

OpenStudy (anonymous):

not trying to get the answer out of you or anything I just learned this concept today so i'm really not as stupid as i seem:)

OpenStudy (anonymous):

give me 1 sec...

OpenStudy (anonymous):

hey its all good I understand! we use multiplication, which means solve this 2 * x = 6

OpenStudy (anonymous):

or x = 6/2

OpenStudy (anonymous):

x=3?

OpenStudy (anonymous):

YES you got it! so we take that 3 and we multiply it by the rest of the equation 5 : 7 + 2 = 7 : 9 x3 x3 x3 x3 x3 15:21 + 6 = 21:28 notice how the 6 is in the same place it was like this one: 5:7 + 2 = 7:9 x:x + 6 = x:x and 15 + 6 = 21 21 + 6 = 28 Everything checks out!

OpenStudy (anonymous):

Just one question isn't the 9 times the 3x actually 27???

OpenStudy (anonymous):

and 21 +6 = 27

OpenStudy (anonymous):

yup sorry lol

OpenStudy (anonymous):

Thanks so much for sticking with me and I was wondering if you could help with a couple more??? I knew it was 27 just messing with you:)

OpenStudy (anonymous):

No problem Boyce! and sure I could help you with a couple more

OpenStudy (anonymous):

Thanks, your awesome!! just give me one sec to transfer my answer.:)

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

Ok here is another... The ratio of the length to the width of a rectangle field is 5:4. The perimeter of the field is 324 meters. What are the dimensions?

OpenStudy (anonymous):

P=2l+2w

OpenStudy (anonymous):

thats the perimeter equation

OpenStudy (anonymous):

You got it.:)

OpenStudy (anonymous):

the perimeter is 324 i just don't get how to get the dimensions

OpenStudy (anonymous):

so it just gives you the perimeter and thats it? no pics or anything?

OpenStudy (anonymous):

All is gives is that little hint that P=2L +2w

OpenStudy (anonymous):

there has to be more information...

OpenStudy (anonymous):

I know I don't know what to tell you. I am confused too.

OpenStudy (anonymous):

i don't know how it is even possible to get an answer from that

OpenStudy (anonymous):

it could be anything from l = 100 w = 62 to l = 50 w = 112

OpenStudy (anonymous):

is it multiple choice?

OpenStudy (anonymous):

no, but give me one second I am going to look at my booklet

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

maybe L=5x and w=4x do you think that is possible???

OpenStudy (anonymous):

why the x's?

OpenStudy (anonymous):

Cause then There is the formula A=lw that I found in the booklet. so Then i could substitute the L=5x and the w=4x

OpenStudy (anonymous):

does it give you area?

OpenStudy (anonymous):

no, but just give me a sec I may have it:)

OpenStudy (anonymous):

ok sweet :)

OpenStudy (anonymous):

Yes I got it do you want to see???

OpenStudy (anonymous):

YESSS lol

OpenStudy (anonymous):

Ok I have to take in this moment it makes me feel so smart:) R u ready??? lol

OpenStudy (anonymous):

DO IT!! hahaha

OpenStudy (anonymous):

So l=5x and w=4x And keep in mend that P=2L +2w 324=2[5x] +2[4x] 324=10x +8x 324=18x 18=x or x=18 The numbers are 5x and 4x so they are: 5(18)=90 4(18)=72 BOOM!!! lol

OpenStudy (anonymous):

look at you!!!! nice!!!

OpenStudy (anonymous):

Can i do another problem with you that is if i don't figure it out first.he he he???

OpenStudy (anonymous):

hahahaha sure!

OpenStudy (anonymous):

ok doky

OpenStudy (anonymous):

Here is another... a^2-25/a+5=6/3 i need to solve by cross multiplying

OpenStudy (anonymous):

I keep getting 35 for my answer but it won't check

OpenStudy (anonymous):

could you start a new question and mention me in it the more comments we put in the slower my computer gets...

OpenStudy (anonymous):

me too sure